# Probing Non-Thermal Processes in Stellar Flares on AU Mic

### v1

@@ Abstract
The role of non-thermal proton energy transportation during solar and stellar flares is largely unknown; a better understanding of this physical process will allow us to rectify longstanding deficiencies in flare models. One way to detect the presence of non-thermal protons during flares is through the Orrall-Zirker (OZ) effect whereby an enhanced red wing appears in hydrogen emission lines (e.g., Lyman- at 1215.67 Å). We analyze archival Hubble Space Telescope/Cosmic Origins Spectrograph G130M (1060 - 1360 Å) observations of the young M dwarf, AU Mic to search for evidence of OZ effect during the impulsive phase of six stellar flares with erg. While we found non-detections of the OZ effect, we note there is a pronounced blue enhancement in several C II and C III emission lines during one of the high-energy flares. We propose that either filament eruptions or chromospheric evaporation could be the mechanism driving this observed blue enhancement. We compare the far-ultraviolet (FUV) spectra to 1D radiative-hydrodynamic stellar flare models, which are unable to reproduce the blue enhancement and broadening in these cool flare lines. By completing a line-by-line analysis of the FUV spectrum of AU Mic, we provide further constraints on the physical mechanisms producing stellar flares on M dwarfs.
@@ 1 Introduction
Non-thermal electrons cannot account for a photospheric origin of white-light flares because they cannot reach deep enough into the atmosphere to cause this heating (Watanabe et al., 2013). To determine the cause of white-light flares and other responses to deep energy deposition, we seek an accurate picture of the non-thermal proton acceleration process. Constraining the presence of low-energy protons ( MeV) among the accelerated particle population in turn constrains the total energies available to explain the white-light radiation and other flare emission. An empirical re-examination of the evidence for these low-energy protons is timely because there has been significant progress in producing power-law spectra of protons extending down to low energies during reconnection (Yin et al., 2024) and in modeling the warm-target transport of low-energy protons from the corona to the chromosphere (Allred et al., 2020). Furthermore, re-examination of gamma-ray images of large solar flares now suggests the sources of non-thermal electron and proton impact with the chromosphere to be cospatial (Battaglia & Krucker, 2025), thus transforming decades of understanding (Hudson, 2016).
@@
One possible observational signature of non-thermal protons accelerated during stellar flare events is the Orrall-Zirker (OZ) effect (Orrall & Zirker, 1976). The OZ effect occurs when non-thermal protons are accelerated downwards in the corona to the mid-upper chromosphere. In the chromosphere, the non-thermal protons charge exchange with ambient hydrogen atoms, thus producing a population of non-thermal energetic neutral atoms (ENAs). These ENAs subsequently go through spontaneous radiative de-excitation and emit Doppler redshifted photons in the red wings of hydrogen spectral lines (Figure 4 in Kerr et al., 2023). Tentative evidence of the OZ effect has been seen in a single star — AU Microscopii (AU Mic; Woodgate et al., 1992). AU Mic is a Myr (Bell et al., 2015) M-dwarf (M1Ve; ; Keenan & McNeil, 1989; Plavchan et al., 2020; Gallenne et al., 2022). AU Mic is known for its enhanced stellar activity due to its young age, including a high flaring rate (Feinstein et al., 2022; Gilbert et al., 2022; Tristan et al., 2023), short rotation period ( days; Plavchan et al., 2020), and strong magnetic field (Donati et al., 2023). AU Mic has been the target for several flare-driven observational campaigns (Tristan et al., 2023; Paudel et al., 2024; Tristan et al., 2025) and is a standard for modeling M-dwarf flares due to its close proximity ( pc; Gaia Collaboration et al., 2021) and high signal-to-noise observations.
@@
Orrall & Zirker (1976) predicted a signature of non-thermal proton beams accelerated during flares could be observed as an increase in the red wing of Lyman- (Ly; 1215.67 Å). Kerr et al. (2023) defined the expected wavelength range to detect the OZ effect as 5.6 - 17.7 Å redward (1380-4365 km s-1) of Ly (see also Canfield & Chang, 1985). In 1992, Woodgate et al. (1992) presented a tentative detection of the OZ effect on AU Mic through an observed red enhancement in Ly using the Goddard High Resolution Spectrograph on the Hubble Space Telescope (HST). The detection lasted for 3 s at the flare start, which is in agreement with the requirement for a relatively large area of undisturbed atmosphere. To date, this is still the only tentative detection of the OZ effect on another star (Kerr et al., 2023; Robinson et al., 2001). This tentative detection highlights that HST could detect the OZ effect here if a sufficient amount of non-thermal protons are accelerated during a flare. Feinstein et al. (2022) performed a preliminary search in recent HST/Cosmic Origins Spectrograph (COS) observations and did not detect the presence of the OZ effect.
@@
In addition to its activity, AU Mic has been a target of observational campaigns due to the presence of two transiting exoplanets (Plavchan et al., 2020; Martioli et al., 2021) and its debris disk (Liu, 2004; Lawson et al., 2023). AU Mic b is a Neptune-like exoplanet () on a close orbit, with a semi-major axis au. AU Mic c is slightly smaller ( ) and farther than AU Mic b, with a semi-major axis of au (Martioli et al., 2021). Due to the close proximity of both planets, their evolution has the potential to be shaped by AU Mic’s frequent stellar flares (e.g., Feinstein et al., 2022; do Amaral et al., 2025). For example, M-dwarf stars’ frequent and intense flaring may be a source of atmospheric stripping of exoplanets Lammer et al. (2007), with exoplanet atmospheric loss being predicted to be partially driven by M-dwarf flaring activity (Neves-Ribeiro et al., 2022). The potential of later atmospheric recovery is debated (Tilley et al., 2019). On the contrary, frequent flaring may jump-start the formation of organic molecules and prebiotic chemistry necessary for the evolution of life (Dransfield et al., 2024); simulations of superflares on the young Sun suggest they may have been instrumental in forming the molecules essential for life (Airapetian et al., 2016).
@@
Beyond the AU Mic system, a complete understanding of the underlying physics driving magnetic reconnection events and the resulting observable stellar flares will provide a comprehensive understanding of the environments in which most planets evolve. M-dwarf stars such as AU Mic are known for their frequent and intense flaring throughout and especially early in their lives (Feinstein et al., 2020b; Crowley et al., 2024). Stellar CMEs, often accompanied by flares, can potentially be associated with spectral line shifts during a stellar flare, as shown in Ikuta & Shibata (2024). Due to the close-in habitable zones around M-dwarfs, atmospheric erosion of potentially habitable planets can be brought about by stellar CMEs from M-dwarfs more easily than CMEs from other spectral types of stars. It is crucial to understand the frequency and time evolution of stellar CMEs from M-dwarfs in order to determine their impact on atmospheric loss and evolution (Khodachenko et al., 2007; Lammer et al., 2007). The current models of stellar flares are not conclusive or complete enough to give accurate insight into the effects of stellar flares on exoplanet atmospheres. Bridging the current gaps in how flares operate will allow an understanding of exoplanet atmospheric responses.
@@
In this work, we present a reanalysis of three HST/COS visits of AU Mic (Feinstein et al., 2022, 2024) to search for evidence of non-thermal proton beams. We present newly identified flares in the third visit. This work is laid out as follows. We describe the HST observations, data reduction, and emission line identification in Section 2. We identify and characterize new stellar flares in Section 3 and present the first far-ultraviolet (FUV) flare-frequency distribution (FFD) for AU Mic. In Section 4, we present the results of our search for evidence of the OZ effect and quantify an identified blue enhancement in several C II and C III emission lines. In Section 5 we speculate further on the cause of the identified carbon enhancement and discuss the implications of this work on stellar flare models. We conclude in Section 6.
@@ 2 Observations
We observed three visits of AU Mic with HST/COS G130M (Å; Å) under GO 16164 (PI Cauley). Visit 1 was executed on 2021-05-28 from 03:33:31 — 15:50:43, Visit 2 was executed on 2021-09-23 from 15:06:14 — 22:02:08, and Visit 3 was executed on 2022-10-09 from 12:38:55 — 19:33:40. The instrument configuration for all three visits is identical. All visits have a total exposure duration of 2085 seconds for the first orbit and an exposure duration of 2471 seconds for the subsequent orbits.
@@ 3 The FUV Stellar Flares of AU Mic
In this work, we present new stellar flares which occurred during Visit 3. We followed the identification technique described in Feinstein et al. (2022). Feinstein et al. (2022) identified flares in Visits 1 and 2 by identifying outliers in the Si III ( = 1294.55 Å) and C III ( = 1175.95 Å) emission line light curves, presented in Figure 1. Flares were identified by their well-known time-dependent behavior of a sharp initial increase followed by an exponential decay phase. We identify an additional seven flares in Visit 3; we follow the same naming convention and introduce flares N - T.
@@
We uniformly calculated the energies and equivalent durations for flares A - T, where flares A - M occurred in visits 1 and 2. We derive the flare energy using the following prescription:
@@
where is the distance to the star, is the start time of the flare defined as the time point the flux increases by from the quiescent flux, is the stop time of the flare defined as the time point the flux gets within of the average quiescent flux, , and is the total flux during the defined start and stop times. We adopted (Gaia Collaboration et al., 2021). To calculate the equivalent duration (ED), first described by Gershberg (1972), we used the following prescription:
@@
We calculated the energies and durations of the new flares in the “white light” (broadband) wavelength-integrated flux ( Å), Si III wavelength-integrated line flux, and C III wavelength-integrated line flux. For the white light flux, we masked the Å due to the lower signal-to-noise at these wavelengths. We repeated this process for the entire sample of flares. We note that Flare R (Figure 1) has an extended decay, which could be due to the presence of lower energy flares we are unable to resolve. This event could be influencing our ability to accurately calculate the true and for Flare S as the flux does not return to the pre-flare baseline before Flare S begins; this can be seen in both the white light and C III light curves and flare properties. The measured properties of all flares are presented in Table 2.
3.1 Optical Flares on AU Mic
AU Mic was observed with the Transiting Exoplanet Survey Satellite, (TESS; Ricker et al., 2015), in Sectors 1 (August 2018) and 27 (July 2020). Gilbert et al. (2022) and Tristan et al. (2025) have both presented independent analyses of the TESS flares on AU Mic. We present a reanalysis using a third independent technique.
We downloaded the TESS 2-minute cadence observations through lightkurve (Lightkurve Collaboration et al., 2018). We use the PDC_SAP flux from the SPOC-processed light curves (Jenkins et al., 2010). We identify flares using the convolutional neural network (CNN) model stella (Feinstein et al., 2020a, b). stella assigns each data point a probability that it is (1) or is not (0) part of a flare. We choose this method as the CNN model is able to identify flares without removing rotational photometric variability. We considered all flares with an assigned probability .
We calculated the flare energies and EDs using Equations 1 and 2. We identified 54 flares in Sector 1 and 55 flares in Sector 27, for a total of 109 flares in 48.7 days of observations. This results in an optical flare rate of 2.24 flares day-1. Our sample of TESS flares has erg. Our recovered number of flares is slightly higher than but still consistent with previous works (Gilbert et al., 2022; Tristan et al., 2025). We note that additional TESS observations of AU Mic are expected in Sector 95 (starting August 2025).
3.2 FUV and Optical Flare Frequency Distributions
We compare the FUV and optical flare frequency distributions (FFD) of AU Mic. We derive our FUV FFD using the white light HST, Si III, and C III flare energies derived in this work. We look at the FFD of multiple spectral lines as this may provide insights into how the emission processes may change over the range of observed flares. We derive our optical FFD using observations from the TESS observations. All resulting FFDs are presented in Figure 2. We fit two distributions to describe the FFD. First, we fit a truncated power law of the form
where is the energy, is a normalization factor, is the slope, is the energy at which the function turns over, is an offset factor, and is the flare rate. We fit this form to characterize . We use the non-linear least squares model optimization implemented with scipy.optimize.curve_fit to fit for each variable. We repeat this process for all FFDs. We fit all available energy bins, with the exception of the highest two energy bins in the HST WL and C III FFDs. We choose to ignore these bins as the high-energy regime of the FFD is sparsely sampled due to the limited baseline of the HST observations.
Additionally, we fit a piecewise function of two power-laws to measure the slope, , of the distribution. We choose to use this as the descriptor of the slope, rather than in the truncated power-law, as measuring is a more common practice when describing FFDs (e.g., Ilin et al., 2019, 2021). We defined our piecewise function as:
where is the slope and is the y-intercept. We present the best-fit values for , , and in Table 1. Due to the large uncertainties in for the HST WL and Si III FFDs, we fit these distributions as a single power-law.
FFD | [erg] | ||
TESS | |||
HST | — | ||
WL | — | ||
HST | — | ||
Si III | |||
HST | |||
C III |
We compare the measured TESS FFD to the average FFD of stars of similar ages and to AU Mic presented in Feinstein et al. (2024). We find that is consistent to within of the FFD slope for stars with K and ages Myr (Feinstein et al., 2024). We find that is inconsistent with the measured FFD slope of stars of similar temperatures and age to AU Mic by . This discrepancy could be due to the high SNR observations of AU Mic, which result in a more complete FFD, or could be an indication that AU Mic is particularly active for its age. A third alternative scenario is that the flares of AU Mic are driven by star-planet interactions with its two short-period gas-giant planets (Ilin & Poppenhaeger, 2022; Feinstein et al., 2024).
Loyd et al. (2018) presented an analysis of inactive and active M dwarfs that were observed as part of the Measurements of the Ultraviolet Spectral Characteristics of Low-mass Exoplanetary Systems (MUSCLES) Treasury survey (France et al., 2016). This study included six “inactive” M dwarfs observed with COS/G130M and four “active” M dwarfs STIS/E140M, including AU Mic. Loyd et al. (2018) measured the FUV flux within specific integrated regions and fit a single power-law with a slope of for active M-dwarfs. We note that the inactive M dwarfs had a flare sample size similar to that presented here. The shallower FFD slope measured here could indicate that AU Mic is more active than the cumulative sample of M-dwarfs. More observations of AU Mic in the FUV are required to properly sample the high-energy regime of the FFD.

### v2

# Far-Ultraviolet Emission Line Investigation of Flares on AU Mic

@@ Abstract
The role of non-thermal proton energy transportation during solar and stellar flares is largely unknown; a better understanding of this physical process will allow us to rectify longstanding deficiencies in flare models. One way to detect the presence of non-thermal protons during flares is through the Orrall-Zirker (OZ) effect, proposed by Orrall & Zirker (1976), whereby an enhanced red wing appears in hydrogen emission lines (e.g., Lyman- at 1215.67 Å). We analyze archival Hubble Space Telescope/Cosmic Origins Spectrograph G130M (1060 - 1360 Å) observations of the young M dwarf, AU Mic to search for evidence of OZ effect during the impulsive phase of six stellar flares with erg. While we found non-detections of the OZ effect, we note there is a pronounced blue enhancement in several C II and C III emission lines during one of the high-energy flares. We propose that either filament eruptions or chromospheric evaporation could be the mechanism driving this observed blue enhancement. We compare the far-ultraviolet (FUV) spectra to 1D radiative-hydrodynamic stellar flare models, which are unable to reproduce the blue enhancement and broadening in these cool flare lines. By completing a line-by-line analysis of the FUV spectrum of AU Mic, we provide further constraints on the physical mechanisms producing stellar flares on M dwarfs.
@@ 1 Introduction
The origin of white light flares is still unknown (Watanabe et al., 2017). Non-thermal electrons alone cannot account for a photospheric origin of white-light flares because they cannot penetrate deep enough into the atmosphere to cause the required heating (Watanabe et al., 2013). Chromospheric evaporation occurs when non-thermal electrons stream along magnetic field lines and impact the chromosphere. Once these non-thermal electrons arrive at the chromosphere, they deposit their energy into the plasma, heating the plasma and increasing pressure. The heated plasma is accelerated upwards, resulting in chromospheric evaporation (Fisher et al., 1985a, b, c).
Observational evidence of chromospheric evaporation includes blue-shifted coronal spectral lines during the impulsive phase of the flare (Milligan & Dennis, 2009; Doschek et al., 1980; Antonucci et al., 1984; Shibata et al., 1995; Graham & Cauzzi, 2015; Polito et al., 2016) and the empirical relation between impulsive and gradual phase emissions known as the Neupert effect (Neupert, 1968; Dennis & Zarro, 1993; Veronig, 2003). There are generally thought to be two regimes of chromospheric evaporation: gentle and explosive, whereby the dividing line is a complicated function of the injected electron beam properties; (flux, low-energy cutoff, and power-law index Fisher, 1989; Reep et al., 2015). Solar models of chromospheric evaporation tend to favor heating driven by electron beams rather than thermal conduction (Polito et al., 2016).
In stellar flares, blueshifts have also been observed in hot coronal flare lines (Argiroffi et al., 2019; Inoue et al., 2024), but occasionally they also appear in cooler chromospheric lines such as H (Fuhrmeister et al., 2008; Honda et al., 2018; Notsu et al., 2024; Lu et al., 2025; Kajikiya et al., 2025). Blueshifts are typically modeled as filament eruptions or coronal mass ejections in stellar flares. Radiative-hydrodynamic models of chromospheric evaporation have been able to produce blueshifts in cool lines (Allred et al., 2005), but the evaporations are too short lived and too slow to explain most of the stellar data. More recent radiative-hydrodynamic models suggest that changes in the electron beam can drive deeper material from the chromosphere into the transition region and corona (Kerr et al., 2024; Polito et al., 2023).
@@
To determine the cause of white-light flares and other responses to deep energy deposition, we seek also an accurate picture of the non-thermal proton acceleration process. Constraining the presence of low-energy protons among the accelerated particle population in turn constrains the total energies available to explain the white-light radiation and other flare emission. An empirical re-examination of the evidence for these low-energy protons is timely because there has been significant progress in producing power-law spectra of protons extending down to low energies during reconnection (Yin et al., 2024) and in modeling the warm-target transport of low-energy protons from the corona to the chromosphere (Allred et al., 2020). Furthermore, re-examination of gamma-ray images of large solar flares now suggests the sources of non-thermal electron and proton impact with the chromosphere to be cospatial (Battaglia & Krucker, 2025), thus transforming decades of understanding (Hudson, 2016).
@@
One possible observational signature of non-thermal protons accelerated during stellar flare events is the Orrall-Zirker (OZ) effect (Orrall & Zirker, 1976). The OZ effect occurs when non-thermal protons are accelerated downwards in the corona to the mid-upper chromosphere. In the chromosphere, the non-thermal protons charge exchange with ambient hydrogen atoms, thus producing a population of non-thermal energetic neutral atoms (ENAs). These ENAs subsequently go through spontaneous radiative de-excitation and emit Doppler redshifted photons in the red wings of hydrogen spectral lines (Figure 4 in Kerr et al., 2023). Tentative evidence of the OZ effect has been seen in a single star — AU Microscopii (AU Mic; Woodgate et al., 1992). AU Mic is a Myr (Bell et al., 2015) M-dwarf (M1Ve; RR; MM Keenan & McNeil, 1989; Plavchan et al., 2020; Gallenne et al., 2022). AU Mic is known for its enhanced stellar activity due to its young age, including a high flaring rate (Feinstein et al., 2022; Gilbert et al., 2022; Tristan et al., 2023), short rotation period ( days; Plavchan et al., 2020), and strong magnetic field (Donati et al., 2023). AU Mic has been the target for several flare-driven observational campaigns (Tristan et al., 2023; Paudel et al., 2024; Tristan et al., 2025) and is a standard for modeling M-dwarf flares due to its close proximity ( pc; Gaia Collaboration et al., 2021) and high signal-to-noise observations.
@@
Orrall & Zirker (1976) predicted a signature of non-thermal proton beams accelerated during flares could be observed as an increase in the red wing of Lyman- (Ly; 1215.67 Å). Kerr et al. (2023) defined the expected wavelength range to detect the OZ effect as 5.6 - 17.7 Å redward (1380-4365 km s-1) of Ly (see also Canfield & Chang, 1985). The OZ effect would be detectable for low energy non-thermal protons with energies ranging from 0.01 to 1 MeV as discussed in Kerr et al. (2023) and Canfield & Chang (1985). Woodgate et al. (1992) presented a tentative detection of the OZ effect on AU Mic through an observed red enhancement in Ly using the Goddard High Resolution Spectrograph on the Hubble Space Telescope (HST). The detection lasted for 3 s at the flare start, which is in agreement with the requirement for a relatively large area of undisturbed atmosphere. To date, this is still the only tentative detection of the OZ effect on another star (Kerr et al., 2023; Robinson et al., 2001). This tentative detection highlights that HST could detect the OZ effect here if a sufficient amount of non-thermal protons are accelerated during a flare ( erg s-1 cm-2, as modeled in Kerr et al. (2023)). Feinstein et al. (2022) performed a preliminary search in recent HST/Cosmic Origins Spectrograph (COS) observations and did not detect the presence of the OZ effect.
@@
In addition to its activity, AU Mic has been a target of observational campaigns due to the presence of two transiting exoplanets (Plavchan et al., 2020; Martioli et al., 2021) and its debris disk (Liu, 2004; Lawson et al., 2023). AU Mic b is a Neptune-like exoplanet (R R⊕) on a close orbit, with a semi-major axis a au. AU Mic c is slightly smaller (R R⊕) and farther than AU Mic b, with a semi-major axis of au (Martioli et al., 2021). Due to the close proximity of both planets, their evolution has the potential to be shaped by AU Mic’s frequent stellar flares (e.g., Feinstein et al., 2022; do Amaral et al., 2025). For example, M-dwarf stars’ frequent and intense flaring may be a source of atmospheric stripping of exoplanets (Lammer et al., 2007), with exoplanet atmospheric loss being predicted to be partially driven by M-dwarf flaring activity (Neves-Ribeiro et al., 2022). The potential of later atmospheric recovery is debated (Tilley et al., 2019). On the contrary, frequent flaring may jump-start the formation of organic molecules and prebiotic chemistry necessary for the evolution of life (Dransfield et al., 2024); simulations of superflares on the young Sun suggest they may have been instrumental in forming the molecules essential for life (Airapetian et al., 2016).
@@
Beyond the AU Mic system, a complete understanding of the underlying physics driving magnetic reconnection events and the resulting observable stellar flares will provide a comprehensive understanding of the environments in which most planets evolve. M-dwarf stars such as AU Mic are known for their frequent and intense flaring throughout and especially early in their lives (Feinstein et al., 2020; Crowley et al., 2024). Stellar CMEs, often accompanied by flares, can potentially be associated with spectral line shifts during a stellar flare, as shown in Ikuta & Shibata (2024). Due to the close-in habitable zones around M-dwarfs, atmospheric erosion of potentially habitable planets can be brought about by stellar CMEs from M-dwarfs more easily than CMEs from other spectral types of stars. It is crucial to understand the frequency and time evolution of stellar CMEs from M-dwarfs in order to determine their impact on atmospheric loss and evolution (Khodachenko et al., 2007; Lammer et al., 2007). The current models of stellar flares are not conclusive or complete enough to give accurate insight into the effects of stellar flares on exoplanet atmospheres. Bridging the current gaps in how flares operate will allow an understanding of exoplanet atmospheric responses. The models used in Chen et al. (2021), where the effects of flares on rocky planets are explored, only use photometric information, so there is no spectroscopic component to these models. By neglecting a spectroscopic component, the full impact of the effects of flares on rocky exoplanets are overlooked (Kowalski et al., 2025). We do not have many exoplanet spectra that show the direct response to flare irradiation to bridge this gap; thus the models of how flares affect exoplanet atmospheres do not leverage the precision that these spectroscopic observations offer.
@@
In this work, we present a reanalysis of three HST/COS visits of AU Mic (Feinstein et al., 2022, 2024) to search for evidence of non-thermal proton beams. We present newly identified flares in the third visit. This work is laid out as follows. We describe the HST observations, data reduction, and emission line identification in Section 2. We identify and characterize new stellar flares in Section 3. In Section 4, we present the results of our search for evidence of the OZ effect and quantify an identified blue enhancement in several C II and C III emission lines. In Section 5 we speculate further on the cause of the identified carbon enhancement and discuss the implications of this work on stellar flare models. We conclude in Section 6.
@@ 2 Observations
We observed three visits of AU Mic with HST/COS G130M (Å; Å, Resolving Power, R= ) under GO 16164 (PI Cauley). Visit 1 was executed on 2021-05-28 from 03:33:31 — 15:50:43, Visit 2 was executed on 2021-09-23 from 15:06:14 — 22:02:08, and Visit 3 was executed on 2022-10-09 from 12:38:55 — 19:33:40. The instrument configuration for all three visits is identical. All visits have a total exposure duration of 2085 seconds for the first orbit and an exposure duration of 2471 seconds for the subsequent orbits.
@@ 3 The FUV Stellar Flares of AU Mic
Understanding the time evolution, energy, and rate of flaring for AU Mic is important in understanding specific physical phenomena, such as the OZ effect, observed during flares. In this work, we present new stellar flares which occurred during Visit 3. We followed the identification technique described in Feinstein et al. (2022). Feinstein et al. (2022) identified flares in Visits 1 and 2 by identifying outliers in the Si III ( = 1294.55 Å) and C III ( = 1175.64 Å) emission line light curves, presented in Figure 1. Flares were identified by their well-known time-dependent behavior of a sharp initial increase followed by an exponential decay phase. We identify an additional seven flares in Visit 3; we follow the same naming convention and introduce flares N - T.
@@
We uniformly calculated the energies and equivalent durations for flares A - T, where flares A - M occurred in visits 1 and 2. We derive the flare energy (E) using the following prescription:
@@
where is the distance to the star, is the start time of the flare defined as the time point the flux increases by 1- from the quiescent flux, is the stop time of the flare defined as the time point the flux gets within 1- of the average quiescent flux, , and is the total flux during the defined start and stop times. The 1- fluxes above , are defined as + erg s-1 cm-2 for the white-light flux, + erg s-1 cm-2 for the Si III flux, and + erg s-1 cm-2 for the C III flux. We adopted 0.002 pc (Gaia Collaboration et al., 2021). To calculate the equivalent duration (ED), first described by Gershberg (1972), we used the following prescription:
@@
We calculated the energies and durations of the new flares in the “white light” (broadband) wavelength-integrated flux ( Å), Si III wavelength-integrated line flux ( Å), and C III wavelength-integrated line flux ( Å). For the white light flux, we masked the Å due to the lower signal-to-noise at these wavelengths. We repeated this process for the entire sample of flares. We note that Flare R (Figure 1) has an extended decay, which could be due to the presence of lower energy flares we are unable to resolve. This event could be influencing our ability to accurately calculate the true and for Flare S as the flux does not return to the pre-flare baseline before Flare S begins; this can be seen in both the white light and C III light curves and flare properties. The measured properties of all flares are presented in Table 1.